# Initial-Condition-Robust Inference in Autoregressive Models

### v1

@@ 1.  Introduction
We consider an AR model with an AR parameter that may take values close or equal to one. Existing CIs in the literature assume that the initial condition is stationary, zero, or fixed, e.g., see [14], [7], [1], [12], [9], [13], and [4, AG14 hereafter]. If the initial condition assumption is not satisfied, then these CIs do not have correct asymptotic coverage probabilities and their finite sample coverage probabilities can be poor. For example, simulations reported below show that the nominal 95% AG14 CI has finite sample coverage probabilities ranging from 24.1% to 93.5% across 50 cases with highly variable initial conditions and different types of conditional heteroskedasticity of the error. A quarter of these CPs are 79.0% or less.
To circumvent the sensitivity of existing CIs to the initial condition, this paper introduces a new initial-condition-robust (ICR) CI whose finite sample CPs do not depend on the initial condition. These CIs are easy to compute and do not depend on any tuning parameters. We show that the ICR CI has correct asymptotic size in a uniform sense under conditions that allow for arbitrary initial condition and conditional heteroskedasticity. Simulations show that the finite sample CPs of the nominal 95% ICR CI are quite good. They lie between 93.5% and 95.0% in the cases considered.
@@
In scenarios where existing CIs are asymptotically valid (i.e., the initial condition is stationary or fixed), the ICR CI pays a price in terms of its length compared to existing CIs. This occurs because the ICR LS estimator has a higher variance due to the additional regressor that is included in the LS regression. However, simulations show that the effect is fairly small. Across 50 scenarios with stationary or fixed initial conditions and i.i.d., GARCH, or ARCH errors, the ratio of the expected length of the ICR CI compared to that of the AG14 CI was found to vary between 1.00 and 1.11. On average, the ICR CIs are 3.5% longer than the AG14 CIs in these scenarios.
@@ 2.  Model and ICR Confidence Interval
where $\rho\in[-1+\varepsilon,1]$ for some $0<\varepsilon<2$ and $\{U_{i}:i=1,2,\dots\}$ are stationary and ergodic under the distribution $F$, with conditional mean 0 given a $\sigma$-field $\mathcal{G}_{i-1}$ for which $U_{j}\in\mathcal{G}_{i}$ for all $j\leq i$, conditional variance $\sigma_{i}^{2}=E_{F}(U_{i}^{2}|\mathcal{G}_{i-1})$, and unconditional variance $\sigma_{U}^{2}\in(0,\infty)$. A detailed description of the parameter space $\Lambda$ is provided in Section 5.1.
@@
$Y_{i}$ $=\mu+\rho^{i-1}\cdot\rho\cdot Y_{0}^{*}+\sum_{j=1}^{n}\rho^{i-j}U_{j}\quad\text{for }i=1,\dots,n.$ (2)
@@ 4.  Monte Carlo Simulations
We compare the coverage probabilities (CPs) and average lengths (ALs) of the ICR and AG14 CIs and report the absolute median biases (ABMs) of the ICR MUE using Monte Carlo simulations. In Section D of the Supplemental Material, we compare the performance of the ICR and [13, Mik07 hereafter] CIs. We focus on the nominal 95% equal-tailed two-sided CIs. For the AG14 CI, we follow the paper’s constructions of the test statistic and critical values. To calculate the MUE for $\rho$, we follow the procedure described in Section 3 and use $\widetilde{\rho}_{n}$ as the MUE of $\rho$ when $\widetilde{\rho}_{n}$ is a singleton. When $\widetilde{\rho}_{n}^{\text{up}}\neq\widetilde{\rho}_{n}^{\text{low}}$, we take $\widetilde{\rho}_{n}^{\text{up}}$ as the MUE for $\rho$ following [5, Section 5.1].
@@
Tables 2 and 3 report the CPs of AG14 and ICR CIs, respectively. The CPs of the AG14 CI in Table 2 are close to the nominal level of 95% when the initial condition is fixed at zero or stationary, but fall below 95% when the initial condition is more variable, namely in the scaled $n$ and explosive $Y_{0}^{*}$ cases. In particular, the AG14 CI is robust to non-i.i.d. innovations when the initial condition is fixed or stationary. On the other hand, for scaled $n$ and explosive $Y_{0}^{*}$ cases, the AG14 CPs lie in the interval [24.1, 93.5] with roughly a quarter of the CPs being less than 79.0. Table 3 shows that the ICR CI achieves CPs close to 95% in all of the scenarios considered and for any initial conditions.
Table 4 reports the ratios of the ALs of the nominal 95% ICR CI to the AG14 CI. We include only cases with fixed or stationary initial conditions because AG14 CIs are not robust to more variable initial conditions, whereas ICR CIs are. The ratios are between 1.00 and 1.11. On average, the ICR CIs are 3.5% longer than the AG14 CIs over the fifty cases considered. Thus, the price the ICR CI pays in terms of AL to gain robustness against the distribution of the initial condition is fairly small.
Table 2: Coverage probabilities ($\times 100$) of nominal 95% AG14 CIs
@@
Table 3: Coverage probabilities ($\times 100$) of nominal 95% ICR CIs
@@
Table 4: Ratios of average lengths of nominal 95% ICR to AG14 CIs
@@
Table 5 presents the ALs of ICR CIs. These are substantially decreasing in $\rho$, as expected, longer for ARCH than i.i.d., and increasing from GARCH1 to GARCH2 to GARCH3.
Table 5: Average lengths of nominal 95% ICR CIs
@@
Table 6: Absolute Median Biases of ICR MUEs
@@ Appendix A Computation of Critical Values
$W(r_{i})=W(r_{i-1})+Z_{i-1}/N^{1/2},\ \ r_{i}=i/N,\ i=1,...,N,$ (14)
where $Z_{i-1}$ are independent standard normal random variables. Next, we compute $I_{h}(r_{j})$, $j=1,\dots,N$, using (10) via a numerical approximation:
@@
Then, we use (10) to compute $I_{f,h}(r_{j})$, $j=1,\dots,N$. Given $I_{f,h}(r_{j})$ and $W(r_{j})$, we evaluate $J_{h}$ in (9) via numerical approximation. Finally, the simulated estimate of $c_{h}(\alpha)$ is obtained as the $\alpha$-th quantile of the empirical distribution of the $B$ realizations of $J_{h}$.
@@ C.1.  Technical Lemmas
When $n(1-\rho_{n})\to h\in(0,\infty)$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, the following results hold jointly
@@
When $n(1-\rho_{n})\to h=0$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, the following results hold jointly
@@
When $n(1-\rho_{n})\to h\in(0,\infty)$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, the following results hold jointly
@@
When $n(1-\rho_{n})\to h=\infty$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, the following results hold jointly
@@
When $n(1-\rho_{n})\to h=\infty$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, the following results hold jointly, we have
@@ C.2.1 Local-to-Unity Case
Since $P(\int_{0}^{1}I_{f,h}^{2}(r)dr=0)=0$, we have $P(\det(B_{n})=0)\to 0$ as $n\to\infty$, which implies that $B_{n}$ is invertible with probability approaching one. Combining it with the fact that $B_{n}=O_{p}(1)$, we have $B_{n}^{-1}=O_{p}(1)$.
@@
$K_{n}$ $=\begin{pmatrix}n^{-2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}&n^{-2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}^{2}&n^{-5/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{3}\\ n^{-2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}^{2}&n^{-2}\sum_{i=2}^{n}\rho_{n}^{2i-2}\widetilde{U}_{i-1}^{2}&n^{-5/2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}^{3}\\ n^{-5/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{3}&n^{-5/2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}^{3}&n^{-3}\sum_{i=2}^{n}\widetilde{U}_{j-1}^{4}\end{pmatrix}$. (28)
@@
$+n^{-2}\frac{1}{(nQ-S^{2})^{2}}\sum_{i=1}^{n}\left(\sum_{k=2}^{n}\left(Q-S(\rho_{n}^{k-1}+\rho_{n}^{i-1})+n\rho^{i+k-2}\right)\widetilde{U}_{k-1}\right)^{2}\widehat{U}_{i}^{2}+o_{p}(1)$
@@ C.3.  Proofs of Lemmas
The proof is thus completed. ∎
@@ Appendix D Additional Simulation Results
We provide additional simulation results in this section. Table SM.1 reports the CPs of [13, Mik07 hereafter] CIs. [^4] The Mik07 CI differs from the [4] CI in that there is no initial-condition component in the asymptotic distribution of the test statistic, and a homoskedastic variance estimator is used. The CPs of Mik07 CIs are close to the nominal level when the $U_{i}$ sequences are i.i.d. and $Y_{0}^{*}$ is either fixed or stationary. When $Y_{0}^{*}$ is drawn from scaled $n$ or explosive distributions, or when the $U_{i}$ sequences are non-i.i.d., the CPs of Mik07 CIs fall below 95%.
Table SM.1: Coverage probabilities ($\times 100$) of nominal 95% Mik07 CIs
@@
Table SM.2 reports the ratios of the ALs of ICR to Mik07 CIs when the $U_{i}$ sequences are i.i.d. and $Y_{0}^{*}$ is fixed or stationary. The ratios are between 1 and 1.12, suggesting that the ALs of ICR CIs are similar to those of Mik07 CIs when the true data generating process is such that the Mik07 inference method is valid. Tables SM.3 and SM.4 report the ALs of [4, AG14 hereafter] and Mik07 CIs, respectively. The results are consistent with the conclusions drawn from Tables 4 and SM.2.
Table SM.2: Ratios of average lengths of nominal 95% ICR to Mik07 CIs
@@
Table SM.3: Average lengths of nominal 95% AG14 CIs
@@
Table SM.4: Average lengths of nominal 95% Mik07 CIs

### v3

@@ 1.  Introduction
We consider an AR model with an AR parameter that may take values close or equal to one. Existing CIs in the literature for the AR parameter assume that the initial condition is stationary, zero, or fixed, e.g., see [15], [7], [1], [12], [9], [14], and [4, AG14 hereafter]. If the initial-condition assumption is not satisfied, then these CIs do not have correct asymptotic coverage probabilities and their finite sample coverage probabilities can be poor. For example, simulations reported below show that the nominal 95% AG14 CI has finite sample coverage probabilities ranging from 24.1% to 93.5% across 50 cases with highly variable initial conditions and different types of conditional heteroskedasticity of the error. A quarter of these CPs are 79.0% or less.
To circumvent the sensitivity of existing CIs to the initial condition, this paper introduces a new initial-condition-robust (ICR) CI whose finite sample CPs do not depend on the initial condition. These CIs are easy to compute and do not depend on any tuning parameters. We show that the ICR CI has correct asymptotic size in a uniform sense under conditions that allow for an arbitrary initial condition and conditional heteroskedasticity. Simulations show that the finite sample CPs of the nominal 95% ICR CI are quite good. They lie between 93.5% and 95.0% in the cases considered.
@@
In scenarios where existing CIs are asymptotically valid (i.e., the initial condition is stationary or fixed), the ICR CI pays a price in terms of its length compared to existing CIs. This occurs because the ICR LS estimator has a higher variance due to the additional regressor that is included in the LS regression. However, simulations show that the effect is fairly small. Across 50 scenarios with stationary or fixed initial conditions and i.i.d., GARCH, or ARCH errors, the ratio of the expected length of the ICR CI compared to that of the AG14 CI is found to vary between 1.00 and 1.11. On average, the ICR CI is 3.5% longer than the AG14 CI in these scenarios.
@@
The AR model considered here has been applied in the literature to exchange rate, commodity and stock prices, and other economic time series, e.g., see [13].
@@ 2.  Model and ICR Confidence Interval
where the observed data are $\{Y_{0},\dots,Y_{n}\}$, $\rho\in[-1+\varepsilon,1]$ for some $0<\varepsilon<2$, and $\{U_{i}:i=1,2,\dots\}$ are stationary and ergodic under the distribution $F$, with conditional mean 0 given a $\sigma$-field $\mathcal{G}_{i-1}$ for which $U_{j}\in\mathcal{G}_{i}$ for all $j\leq i$, conditional variance $\sigma_{i}^{2}=E_{F}(U_{i}^{2}|\mathcal{G}_{i-1})$, and unconditional variance $\sigma_{U}^{2}\in(0,\infty)$. A detailed description of the parameter space $\Lambda$ is provided in Section 5.1.
@@
$Y_{i}$ $=\mu+\rho^{i-1}\cdot\rho\cdot Y_{0}^{*}+\sum_{j=1}^{i}\rho^{i-j}U_{j}\quad\text{for }i=1,\dots,n.$ (2)
@@ 4.  Monte Carlo Simulations
We compare the coverage probabilities (CPs) and average lengths (ALs) of the ICR and AG14 CIs and report the absolute median biases (ABMs) of the ICR MUE using Monte Carlo simulations. In Section D of the Supplemental Material, we compare the performance of the ICR and [14, Mik07 hereafter] CIs. We focus on the nominal 95% equal-tailed two-sided CIs. For the AG14 CI, we follow the paper’s constructions of the test statistic and critical values. To calculate the MUE for $\rho$, we follow the procedure described in Section 3 and use $\widetilde{\rho}_{n}$ as the MUE of $\rho$ when $\widetilde{\rho}_{n}$ is a singleton. When $\widetilde{\rho}_{n}^{\text{up}}\neq\widetilde{\rho}_{n}^{\text{low}}$, we take $\widetilde{\rho}_{n}^{\text{up}}$ as the MUE for $\rho$ following [5, Section 5.1].
@@
Tables 2 and 3 report the CPs of AG14 and ICR CIs, respectively. The CPs of the AG14 CI in Table 2 are close to the nominal level of 95% when the initial condition is fixed at zero or stationary, but fall below 95% when the initial condition is more variable, namely in the scaled $n$ and explosive $Y_{0}^{*}$ cases. In particular, the AG14 CI is robust to non-i.i.d. innovations when the initial condition is fixed or stationary. On the other hand, for scaled $n$ and explosive $Y_{0}^{*}$ cases, the AG14 CPs lie in the interval [24.1, 93.5] with roughly a quarter of the CPs being less than 79.0. Table 3 shows that the ICR CI achieves CPs close to 95% in all of the scenarios considered and for any initial conditions. Specifically, all ICR CI CPs lie in the interval [93.5, 95.0].
Table 4 reports the ratios of the ALs of the nominal 95% ICR CI to the AG14 CI. We include only cases with fixed or stationary initial conditions because the AG14 CI is not robust to more variable initial conditions, whereas the ICR CI is. The ratios are between 1.00 and 1.11. On average, the ICR CI is 3.5% longer than the AG14 CI over the fifty cases considered. Thus, the price the ICR CI pays in terms of AL to gain robustness against the distribution of the initial condition is fairly small.
Table 2: Coverage probabilities ($\times 100$) of the nominal 95% AG14 CI
@@
Table 3: Coverage probabilities ($\times 100$) of the nominal 95% ICR CI
@@
Table 4: Ratios of the average lengths of the nominal 95% ICR to AG14 CIs
@@
Table 5: Average lengths of the nominal 95% ICR CI
@@
Table 5 presents the ALs of the ICR CI. These are substantially decreasing in $\rho$, as expected, longer for ARCH conditional heteroskedasticity than i.i.d., and increasing from GARCH1 to GARCH2 to GARCH3 conditional heteroskedasticity.
Table 6: Absolute Median Biases of the ICR MUE
@@ Appendix A Computation of Critical Values
$W(r_{i})=W(r_{i-1})+Z_{i-1}/N^{1/2}\ \text{ for }r_{i}=i/N\text{ and }i=1,...,N,$ (14)
where $Z_{i-1}$ are independent standard normal random variables. Next, we compute $I_{h}(r_{j})$ for $j=1,\dots,N$, using (10) via a numerical approximation:
@@
Then, we use (10) to compute $I_{f,h}(r_{j})$ for $j=1,\dots,N$. Given $I_{f,h}(r_{j})$ and $W(r_{j})$, we evaluate $J_{h}$ in (9) via numerical approximation. Finally, the simulated estimate of $c_{h}(\alpha)$ is obtained as the $\alpha$-th quantile of the empirical distribution of the $B$ realizations of $J_{h}$.
@@ C.1.  Technical Lemmas
When $n(1-\rho_{n})\to h\in(0,\infty)$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, the following results hold jointly:
@@
When $n(1-\rho_{n})\to h=0$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, the following results hold jointly:
@@
When $n(1-\rho_{n})\to h\in(0,\infty)$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, the following results hold jointly:
@@
When $n(1-\rho_{n})\to h=\infty$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, the following results hold jointly:
@@
When $n(1-\rho_{n})\to h=\infty$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, the following results hold jointly:
@@ C.2.1 Local-to-Unity Case
Since $P(\int_{0}^{1}I_{f,h}^{2}(r)dr=0)=0$, we have $P(\det(B_{n})=0)\to 0$ as $n\to\infty$, which implies that $B_{n}$ is invertible with probability approaching one. Combining this with the fact that $B_{n}=O_{p}(1)$, we have $B_{n}^{-1}=O_{p}(1)$.
@@
$K_{n}$ $=\begin{pmatrix}n^{-2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}&n^{-2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}^{2}&n^{-5/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{3}\\ n^{-2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}^{2}&n^{-2}\sum_{i=2}^{n}\rho_{n}^{2i-2}\widetilde{U}_{i-1}^{2}&n^{-5/2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}^{3}\\ n^{-5/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{3}&n^{-5/2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}^{3}&n^{-3}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{4}\end{pmatrix}$. (28)
@@
$+n^{-2}\frac{1}{(nQ-S^{2})^{2}}\sum_{i=1}^{n}\left(\sum_{k=2}^{n}\left(Q-S(\rho_{n}^{k-1}+\rho_{n}^{i-1})+n\rho_{n}^{i+k-2}\right)\widetilde{U}_{k-1}\right)^{2}\widehat{U}_{i}^{2}+o_{p}(1)$
@@ C.3.  Proofs of Lemmas
Thus, the proof is completed.. ∎
@@ Appendix D Additional Simulation Results
We provide additional simulation results in this section. Table SM.1 reports the CPs of the [14, Mik07 hereafter] CI. [^4] The Mik07 CI differs from the [4] CI in that there is no initial-condition component in the asymptotic distribution of the test statistic, and a homoskedastic variance estimator is used. The CPs of the Mik07 CI are close to the nominal level when the $U_{i}$ sequences are i.i.d. and $Y_{0}^{*}$ is either fixed or stationary. When $Y_{0}^{*}$ is drawn from scaled $n$ or explosive distributions, or when the $U_{i}$ sequences are non-i.i.d., the CPs of the Mik07 CI fall below 95%.
Table SM.1: Coverage probabilities ($\times 100$) of the nominal 95% Mik07 CI
@@
Table SM.2 reports the ratios of the ALs of ICR to Mik07 CIs when the $U_{i}$ sequences are i.i.d. and $Y_{0}^{*}$ is fixed or stationary. The ratios are between 1 and 1.12, suggesting that the ALs of the ICR CI are similar to those of the Mik07 CI when the true data generating process is such that the Mik07 inference method is valid. Tables SM.3 and SM.4 report the ALs of [4, AG14 hereafter] and the Mik07 CI, respectively. The results are consistent with the conclusions drawn from Tables 4 and SM.2.
Table SM.2: Ratios of the average lengths of the nominal 95% ICR to Mik07 CIs
@@
Table SM.3: Average lengths of the nominal 95% AG14 CI
@@
Table SM.4: Average lengths of the nominal 95% Mik07 CI